# SUBA-Jet: a new Model for Jets in Heavy Ion Collisions

### v1

@@ V.1 Benchmark of the Gunion-Bertsch radiation seed
Figure 14: Top: Radiation spectrum of virtual gluons immediately after their creation, i.e. before the LPM evolution, in our Monte Carlo approach. We display the result for different choices of the mass of the QGP partons, $m_{q}$, as compared to the superficial $\omega^{-1}$ dependence of the Gunion Bertsch radiation formula with an arbitrary normalization (see text for details). Bottom: Ratio of the spectrum for different choices of $m_{q}$ and the Gunion-Bertsch $1/\omega$ expectation.
In Fig. 14 we present the energy distribution of the virtual gluons immediately after the generating inelastic collision, corresponding to the integration of eq. (29) over $\mathrm{d}^{2}k_{{\perp}}\,\mathrm{d}^{2}q_{{\perp}}$. We display the result of our Monte Carlo sampling for two cases: For a mass of the QGP parton of $m_{\text{q}}=\infty$ (solid blue line) and of $m_{\text{q}}=0$ (dashed orange line). These curves are compared to the generic spectrum of the GB approach, which falls off $\propto\omega^{-1}$. At intermediate $\omega$, sufficiently large that masses and the infrared cutoff $\mu$ do not play a role and sufficient small that the energy/momentum conservation can be neglected, our sampled spectrum agrees with that from the GB approach. The finite gluon mass enters our cross section formula via $\tilde{m}_{g}^{2}(x)=(1-x)m_{g}^{2}+x^{2}m_{Q}^{2}$. 202020 In particular, $m_{g}$ is taken as $m_{g}^{\rm reg}$ for $\omega\gtrsim T$, which is quite small, resulting in $\tilde{m}_{g}^{2}(x)\approx x^{2}m_{Q}^{2}$. For the $\mu$ dependence we refer to eq. (105) for $m_{q}=0$ and to eq. (131) for $m_{q}=\infty$. The choice of $m_{q}$ has an effect on the overall normalization of the curves but less on the shape
@@
The $k_{\perp}$ distribution of the virtual gluons directly after their creation, is shown in Fig. 15. Here we observe that for $k_{{\perp}}<1$ GeV the choice of $m_{q}$ changes only the normalization of the distribution, while for the large $k_{{\perp}}$ energy/momentum conservation modifies the shape differently. From eqs. (79) and (122), derived in appendix B and C, one realizes that $k_{{\perp}}^{\rm max}$ can be as large as the jet energy $E_{Q}$ for $m_{q}=\infty$ case while it saturates at $\sqrt{s}\sim\sqrt{E_{Q}e_{q}}$ for $m_{q}=0$, with $e_{q}$ being the energy of the QGP parton before scattering. In the former case, one thus explores a larger range of $k_{{\perp}}$ values, with $(1/k_{{\perp}})\mathrm{d}N/\mathrm{d}k_{{\perp}}\sim k_{{\perp}}^{-4}$, as can clearly be seen. In the later case, the $k_{{\perp}}$ range is smaller, following the thermal distribution of partons in the QGP, and no algebraic law can be observed. We display as well a reference spectrum $\propto 1/k_{{\perp}}^{4}$, the analytical result of the GB cross section for $q_{{\perp}}\ll k_{{\perp}}$, for the later discussion of formed gluons.
@@
In Figs. 17 and 17, we present the double differential spectrum $\mathrm{d}^{2}N/(\mathrm{d}k_{z}\,k_{{\perp}}\,\mathrm{d}k_{{\perp}})$ for both choices of $m_{q}$. We see that for $m_{q}\to\infty$ a considerable fraction of gluons has even negative $k_{z}$ momenta. We see also clearly how the energy/momentum conservation limits the available phase space in the $m_{q}=0$ case. In the region, which in both cases is kinematically allowed, the distributions are rather similar.
@@ V.2 Reproduction of the BDMPS-Z and GLV limits
In Fig. 18 and  19 gluon energy spectra per unit length $(1/L)\,\mathrm{d}N/\mathrm{d}\omega$ are displayed. In Fig. 18 we consider a medium of length $L=4$ fm and an initial projectile quark energy of $E=100$ TeV to numerically approximate the $E\to\infty$ limit, whereas in Fig. 19 the medium length is 8 fm and the initial projectile quark energy is $E=100$ GeV. For these parameters, we have, according to eq. (39), $\omega_{c}\approx 96$ GeV and $385$ GeV, respectively. The BDMPS-Z results are calculated with eq. (34) and not in the limit $L\to\infty$ (eq (37)).
In Fig. 18 we see that for intermediate energies $\omega<\omega_{c}$, the spectrum falls off roughly as $\omega^{-3/2}$, as expected by BDMPS-Z. This demonstrates that our Monte Carlo algorithm is able to reproduce the interference of the amplitudes. For $\omega\gg\omega_{c}$, the spectrum falls off like $\omega^{-2}$, as predicted in the GLV approach and not like $\omega^{-3}$ in the BDMPS-Z approach above $\omega_{c}$). For this kinematic condition ($E=100$ TeV and $L=4$ fm) around $300$ virtual gluons are produced but only a bit more than $1\%$ become real gluons.
@@
In Fig. 19, the gluon energy spectrum per unit length fulfills as well the BDMPS-Z expectation for intermediate energies, although the initial parton energy is much smaller. The path length is larger than in Fig. 18. We see a suppression of the spectrum at larger $\omega$ resembling the GLV expectation even though we have $\omega_{c}\gg E$. This suppression is visible already in the distribution for the virtual gluons, before the LPM evolution, where both energy conservation constraint and the $1-x$ cofactor in the GB formula suppress the radiation at large $\omega$ with respect to a $1/\omega$ spectrum.
@@
Figs. 21 and 21 show the double differential spectrum $\mathrm{d}N^{2}/(\mathrm{d}N_{s}\,\mathrm{d}\omega)$ for $L=4$ fm, $E=100$ TeV and $L=8$ fm, $E=100$ GeV, respectively. The number of elastic scatterings $N_{s}$ suffered by the gluon is proportional to the formation time, $N_{s}\sim t_{f}/\lambda_{\rm el}^{g}$. The formation time in the LPM (BDMPS-Z) is expected to behave as $t_{f}\sim\sqrt{\omega}$. The red lines on Figs. 21 and 21 indicate the average $N_{s}$ as a function of the gluon energy $\omega$, and approximates a linear dependence for intermediate energies as expected on a log-log plot. For small $\omega$, the early decrease of $\langle N_{s}\rangle$ with $\omega$ reflects the opening of the accessible phase space with $\omega$ which makes these collisions more efficient in ‘‘dephasing’’ the virtual gluon. For $\omega>\omega_{c}$ we see that the average $N_{s}$ decreases. This is a clear sign that this $\omega$-range corresponds to another regime, the GLV regime. In this limit there is (at least) one hard scattering which can be accompanied by soft collisions. Therefore, in our MC, one does not strictly reach $N_{s}=1$ even for $\omega\gg\omega_{c}$. One should notice the large fluctuations around the average $\langle N_{s}\rangle$, which correspond to fluctuations in the formation length.
@@ V.3 Testing the role of the prescriptions in the elastic collisions
Fig. 24 shows the transverse momentum spectrum of the produced gluons for the different prescriptions of the elastic scattering of the virtual gluons. The color coding is the same as in ig. 22. In addition, we add the distribution of the transverse gluon momentum from the initial GB seed. We see, as expected from eq. (33), a large reduction with regard to the GB seed for small $k_{\perp}$, indicating that these gluons have a large formation time and are produced after multiple scatterings — and thus according to the BDMPS-Z scheme.
We see also a significant difference between the different prescriptions for low $k_{\perp}$. This is not surprising because the different prescriptions of the energy momentum conservation in the elastic scattering of virtual gluons yield different values of $k_{\perp}$ if $k_{\perp}$ is small. In particular we observe an excess of soft gluons in the energy-reduction case whereas the $k^{+}$ conservation prescription shifts the gluons to larger values of $k_{\perp}$ and depopulates the low $k_{\perp}$ region. Only when $k_{\perp}$ is of the order $1$ GeV or larger, the three prescriptions yield similar results, a spectrum falling off approximately like $1/k_{\perp}^{4}$.
For $k_{\perp}>1$ GeV, the $k_{\perp}$ spectrum of radiated gluons gradually merges with the one for the virtual gluons, which indicates that very few elastic collisions (if at all) are necessary to form these gluons, because the initial virtual gluons are already sampled with a large $k_{{\perp}}$, and therefore $N_{s}$ is small. This is the case for $k_{\perp}^{2}\gtrsim\omega\varphi_{c}/L$. This confirms that our model is able to interpolate between the low and high $k_{\perp}$ regimes. [^22] The BDMPS-Z spectrum is approximately given by $\mathrm{d}N/\mathrm{d}\omega\sim\omega^{-3/2}$. At the time the gluon is formed, it has $k_{\perp}^{2}=\hat{q}\times\sqrt{\omega/\hat{q}}=\sqrt{\omega\hat{q}}$, meaning that $\omega=k_{\perp}^{4}/\hat{q}$. Then, one has $\frac{1}{k_{\perp}}\mathrm{d}N/\mathrm{d}k_{\perp}=\frac{1}{k_{\perp}}\mathrm{d}N/\mathrm{d}\omega\times\mathrm{d}\omega/\mathrm{d}k_{\perp}\propto 1/k_{\perp}^{4}$, similar to the GB distribution. Hence, getting $\frac{1}{k_{\perp}}\mathrm{d}N/\mathrm{d}k_{\perp}\propto 1/k_{\perp}^{4}$ is not specific to the GLV regime, but observing no reduction with regard to the GB seed on Fig. 24 for $k_{\perp}\gtrsim 4$ GeV$/c$ is.
@@
The distribution of the number of elastic collisions, $N_{s}$, which the virtual gluon has to suffer before it gets formed, is shown in Fig. 25, as well for the three different prescriptions in the elastic scatterings. Here, all approaches give very similar results for $N_{s}>10$, although the distribution for the energy-reduction option appears to be slightly depleted. However, for values of $N_{s}$ around 5, typically associated with small $\omega$ values through $N_{s}\sim t_{f}\sim\sqrt{\omega}$, the energy reduction prescription gives a larger contribution. This has two origins. On the one side, the energy loss in the individual elastic collisions (see Fig. 23) drives some gluons to the soft sector for which the formation time (and so the associated $N_{s}$) is smaller. Because for finite gluon masses the formation time is $\propto\omega/(m_{g}^{2}+k_{{\perp}}^{2})$, the shift of the gluons to smaller $k_{{\perp}}$ values, seen in Fig. 24, does not lower the formation time considerably. For the ‘‘energy reduction’’ scenario we observe therefore an increase of the production for low $N_{S}$. This effect is reinforced by a kinematic effect: The smaller the value of $\omega$ the larger is the probability that elastic collisions of the virtual gluons are kinematically not possible. Then the collisions do not take place and $N_{s}$ is not increased but the phase accumulation continues. This also leads to an enhancement of the entry at $N_{s}=1$, as seen in Fig. 25.
@@ V.4 Testing the role of the phase accumulation
Another quantity, which is approximated in the BDMPS-Z conditions, is the phase accumulation of the virtual gluon due to elastic collisions. Figs. 27, 27, and 28 show the influence of different forms of the phase accumulation of the gluon emission amplitudes on the gluon observables. The form of the phase accumulation cannot be inferred from more fundamental approaches. As discussed in section IV the virtual gluon will be considered as a formed particle once its phase passes the critical value of $\varphi_{c}=6$. We display the results for three different descriptions of this phase accumulation. The results for the phase increment of $\Delta\varphi=\frac{2P\cdot k^{\prime}}{E}\cdot\Delta t$ are shown as a blue solid line, that of a phase increment of $\Delta\varphi=\frac{k_{\perp}^{2}}{\omega}\cdot\Delta t$, corresponding to a vanishing gluon mass, as a dashed orange line and that of a phase increment of $\Delta\varphi=\frac{m_{g}^{2}+k_{\perp}^{2}}{\omega}\cdot\Delta t$, corresponding to the limit $E\to\infty$, as a green dotted line.
@@
For both, the gluon $\omega$ and the $k_{\perp}$ spectra, the high energy limit $\Delta\varphi=\frac{m_{g}^{2}+k_{\perp}^{2}}{\omega}\cdot\Delta t$ presents a reasonable approximation of the more exact formula $\Delta\varphi=\frac{2P\cdot k^{\prime}}{E}\cdot\Delta t$.

### v2

@@ V.1 Benchmark of the Gunion-Bertsch radiation seed
Figure 18: Top: Radiation spectrum of virtual gluons immediately after their creation, i.e. before the LPM evolution, in our Monte Carlo approach. We display the result for different choices of the mass of the QGP partons, $m_{q}$, as compared to the superficial $\omega^{-1}$ dependence of the Gunion Bertsch radiation formula with an arbitrary normalization (see text for details). $N_{g}^{\rm preformed}$ is the number of preformed gluons. Bottom: Ratio of the spectrum for different choices of $m_{q}$ and the Gunion-Bertsch $1/\omega$ expectation.
In Fig. 18 we present the energy distribution of the virtual gluons immediately after the generating inelastic collision, corresponding to the integration of eq. (30) over $\mathrm{d}^{2}k_{{\perp}}\,\mathrm{d}^{2}q_{{\perp}}$. We display the result of our Monte Carlo sampling for two cases: For a mass of the QGP parton of $m_{\text{q}}=\infty$ (solid blue line) and of $m_{\text{q}}=0$ (dashed orange line). These curves are compared to the generic spectrum of the GB approach, which falls off $\propto\omega^{-1}$. At intermediate $\omega$, sufficiently large that masses and the infrared cutoff $\mu$ do not play a role and sufficient small that the energy/momentum conservation can be neglected, our sampled spectrum agrees with that from the GB approach. The finite gluon mass enters our cross section formula via $\tilde{m}_{g}^{2}(x)=(1-x)m_{g}^{2}+x^{2}m_{Q}^{2}$. 222222 In particular, $m_{g}$ is taken as $m_{g}^{\rm reg}$ for $\omega\gtrsim T$, which is quite small, resulting in $\tilde{m}_{g}^{2}(x)\approx x^{2}m_{Q}^{2}$.. For the $\mu$ dependence we refer to eq. (111) for $m_{q}=0$ and to eq. (139) for $m_{q}=\infty$. The choice of $m_{q}$ has an effect on the overall normalization of the curves but less on the shape. At small $\omega$, the shrinking of the phase space $\Phi$, discussed in appendix B – see in particular eqs. (91) and (92) – tames the $\omega^{-1}$ IR divergence found in the GB spectrum and even bends it down towards zero when $\omega\searrow m_{g}$, as previously discussed in Aichelin:2013mra . By properly implementing this phase space boundary, we do not need to resort to some artificial IR regulator.
@@
The $k_{\perp}$ distribution of the virtual gluons directly after their creation, is shown in Fig. 19. Here we observe that for $k_{{\perp}}<1$ GeV the choice of $m_{q}$ changes only the normalization of the distribution, while for the large $k_{{\perp}}$ energy/momentum conservation modifies the shape differently. From eqs. (85) and (128), derived in appendix B and B.4, one realizes that $k_{{\perp}}^{\rm max}$ can be as large as the jet energy $E_{Q}$ for $m_{q}=\infty$ case while it saturates at $\sqrt{s}\sim\sqrt{E_{Q}e_{q}}$ for $m_{q}=0$, with $e_{q}$ being the energy of the QGP parton before scattering. In the former case, one thus explores a larger range of $k_{{\perp}}$ values, with $(1/k_{{\perp}})\mathrm{d}N/\mathrm{d}k_{{\perp}}\sim k_{{\perp}}^{-4}$, as can clearly be seen. In the later case, the $k_{{\perp}}$ range is smaller, following the thermal distribution of partons in the QGP, and no algebraic law can be observed. We display as well a reference spectrum $\propto 1/k_{{\perp}}^{4}$, the analytical result of the GB cross section for $q_{{\perp}}\ll k_{{\perp}}$, for the later discussion of formed gluons.
@@
In Figs. 21 and 21, we present the double differential spectrum $\mathrm{d}^{2}N/(\mathrm{d}k_{z}\,k_{{\perp}}\,\mathrm{d}k_{{\perp}})$ for both choices of $m_{q}$. We see that for $m_{q}\to\infty$ a considerable fraction of gluons has even negative $k_{z}$ momenta. We see also clearly how the energy/momentum conservation limits the available phase space in the $m_{q}=0$ case. In the region, which in both cases is kinematically allowed, the distributions are rather similar.
@@ V.2 Reproduction of the BDMPS-Z and GLV limits
In Figs. 22 and  23 gluon energy spectra per unit length $(1/L)\,\mathrm{d}N/\mathrm{d}\omega$ are displayed. In Fig. 22 we consider a medium of length $L=4$ fm and an initial projectile quark energy of $E=100$ TeV to numerically approximate the $E\to\infty$ limit, whereas in Fig. 23 the medium length is 8 fm and the initial projectile quark energy is $E=100$ GeV. For these parameters, we have, according to eq. (40), $\omega_{c}\approx 96$ GeV and $385$ GeV, respectively. The BDMPS-Z results are calculated with eq. (35) and not in the limit $L\to\infty$ (eq (38)).
In Fig. 22 we see that for intermediate energies $\omega<\omega_{c}$, the spectrum falls off roughly as $\omega^{-3/2}$, as expected by BDMPS-Z. This demonstrates that our Monte Carlo algorithm is able to reproduce the interference of the amplitudes. For $\omega\gg\omega_{c}$, the spectrum falls off like $\omega^{-2}$, as predicted in the GLV approach and not like $\omega^{-3}$ as in the BDMPS-Z approach above $\omega_{c}$). For this kinematic condition ($E=100$ TeV and $L=4$ fm) around $300$ virtual gluons are produced but only a bit more than $1\%$ become real gluons.
@@
In Fig. 23, the gluon energy spectrum per unit length fulfills as well the BDMPS-Z expectation for intermediate energies, although the initial parton energy is much smaller. The path length is larger than in Fig. 22. We see a suppression of the spectrum at larger $\omega$ resembling the GLV expectation even though we have $\omega_{c}\gg E$. This suppression is visible already in the distribution for the virtual gluons, before the LPM evolution, where energy and momentum conservation constraints and the $1-x$ cofactor in the GB formula suppress the radiation at large $\omega$ with respect to a $1/\omega$ spectrum. At low $\omega$, the agreement with the BDMPS-Z spectrum extends down to $\omega$ as small as $\omega\approx 2\,{\rm GeV}$, what should be taken with a grain of salt (see discussion of Fig. 26 ) .
@@
Figs. 25 and 25 show the double differential spectrum $\mathrm{d}N^{2}/(\mathrm{d}N_{s}\,\mathrm{d}\omega)$ for $L=4$ fm, $E=100$ TeV and $L=8$ fm, $E=100$ GeV, respectively. The number of elastic scatterings $N_{s}$ suffered by the gluon is proportional to the formation time, $N_{s}\sim t_{f}/\lambda_{\rm el}^{g}$. The formation time in the LPM (BDMPS-Z) is expected to behave as $t_{f}\sim\sqrt{\omega}$. The red lines on Figs. 25 and 25 indicate the average $N_{s}$ as a function of the gluon energy $\omega$, and approximates a linear dependence for intermediate energies as expected on a log-log plot. For small $\omega$, the early decrease of $\langle N_{s}\rangle$ with $\omega$ reflects the opening of the accessible phase space with $\omega$ which makes these collisions more efficient in ‘‘dephasing’’ the virtual gluon. For $\omega>\omega_{c}$ we see that the average $N_{s}$ decreases. This is a clear sign that this $\omega$-range corresponds to another regime, the GLV regime. In this limit there is (at least) one hard scattering which can be accompanied by soft collisions. Therefore, in our MC, one does not strictly reach $N_{s}=1$ even for $\omega\gg\omega_{c}$. One should notice the large fluctuations around the average $\langle N_{s}\rangle$, which correspond to fluctuations in the formation length.
Figure 26: The gluon $k_{\perp}$ spectrum from our Monte Carlo approach. To match the BDMPS-Z conditions we assume infinitely massive scattering centers and energy conservation for elastic scattering. The gluon phase accumulation is calculated according to $\Delta\varphi=\left(2P_{Q}\cdot k/E_{Q}\right)\Delta t$. The theoretical predictions calculated in the BDMPS-Z harmonic approximation Wiedemann:2000tf is also shown.
In Fig. 26, we compare the $1/k_{\perp}dN_{g}/k_{\perp}$ spectrum from our MC simulation with the equivalent spectrum calculated in the BDMPS-Z harmonic approximation Wiedemann:2000tf , restricting the $k_{z}$ range to $k_{z}>k_{\perp}$ in order to satisfy at best the BDMPS-Z conditions. It is known that the average $\langle k_{\perp}^{2}\rangle$ scales like $Q_{s}^{2}=\hat{q}L$ (see the discussion in Zapp:2011ya as well as in chapter 3 of caucal:tel-03081993 ) and we indeed observe a turning point of both spectra for such a value, with $1/k_{\perp}dN_{g}/k_{\perp}\propto Q_{s}/k_{\perp}^{4}$ for $k_{\perp}^{2}\gtrsim Q_{s}$ Although our MC is in good agreement with the BDMPS-Z calculation in this region, one observes some discrepancy for $k_{\perp}^{2}\lesssim Q_{s}$. One reason for this discrepancy is that the integrated $dN_{g}/k_{\perp}$ spectrum is in fact dominated by low values of $\omega$ and $k_{z}$ – as can be seen from Fig. 39 –, which do not correspond to the BDMPS-Z conditions. A more realistic comparison with the BDMPS-Z calculation would require to select "large" values of $\omega$ (or $k_{z}$) and study both spectra for $k_{\perp}\ll\omega$. We plan to perform such task in a future work, comparing not only to the BDMPS-Z harmonic approximation but also to the full case Isaksen:2023nlr , which requires a numerical calculation in its own. Finally, one may wonder why the $dN/d\omega$ spectrum displayed on Fig. 23 does not reflect, in a way or another, the discrepancy observed on Fig. 26 around $k_{\perp}\approx Q_{s}$, as we have just claimed that $dN/dk_{\perp}$ is mostly sensitive to regions of the double differential spectrum $d^{2}N/dk^{+}dk_{\perp}$ located around $k^{+}=k_{\perp}$. The reason is that eq.(39) results from the integration of $d^{2}N/dk^{+}dk_{\perp}$ over the full $k_{\perp}$ range, implying regions $k_{\perp}\gtrsim k^{+}$ for which the BDMPS-Z condition is not satisfied, while imposing the $k_{\perp}\lesssim k^{+}$ condition – as done to obtain the $k_{\perp}$ spectrum presented in Fig.26 – the integration leads to substantial depletion with respect to eq. (39) for $\omega\lesssim Q_{s}$.
@@ V.3 Testing the role of the prescriptions in the elastic collisions
Fig. 29 shows the transverse momentum spectrum of the produced gluons for the different prescriptions of the elastic scattering of the virtual gluons. The color coding is the same as in Fig. 27. In addition, we add the distribution of the transverse gluon momentum from the initial GB seed. We see, as expected from eq. (34), a large reduction with regard to the GB seed for small $k_{\perp}$, indicating that these gluons have a large formation time and are produced after multiple scatterings — and thus according to the BDMPS-Z scheme.
We see also a significant difference between the different prescriptions for low $k_{\perp}$. This is not surprising because the different prescriptions of the energy momentum conservation in the elastic scattering of virtual gluons yield different values of $k_{\perp}$ if $k_{\perp}$ is small. In particular we observe an excess of soft gluons in the energy-reduction case whereas the $k^{+}$ conservation prescription shifts the gluons to larger values of $k_{\perp}$ and depopulates the low $k_{\perp}$ region. Only when $k_{\perp}$ is of the order $1$ GeV or larger, the three prescriptions yield similar results, a spectrum falling off approximately like $1/k_{\perp}^{4}$. Such excess at low $k_{\perp}$ in the energy reduction case can be understood as follows: Most of these virtual gluons are in fact soft, implying that they do not possess enough energy to allow for the recoil of the medium parton (for the preselected $l_{T}$). The corresponding collisions are therefore vetoed, implying that the virtual gluon effectively propagates through a more dilute medium as for the energy conservation case, implying that they are less impacted by the LPM effect.
For $k_{\perp}>1$ GeV, the $k_{\perp}$ spectrum of radiated gluons gradually merges with the one for the virtual gluons, which indicates that very few elastic collisions (if at all) are necessary to form these gluons, because the initial virtual gluons are already sampled with a large $k_{{\perp}}$, and therefore $N_{s}$ is small. This is the case for $k_{\perp}^{2}\gtrsim\omega\varphi_{c}/L$. This confirms that our model is able to reproduce the shape of the large-$k_{T}$ (GLV) regime.
@@
The distribution of the number of elastic collisions, $N_{s}$, which the virtual gluon has to suffer before it gets formed, is shown in Fig. 30, as well for the three different prescriptions in the elastic scatterings. Here, all approaches give very similar results for $N_{s}>10$, although the distribution for the energy-reduction option appears to be slightly depleted. However, for values of $N_{s}$ around 5, typically associated with small $\omega$ values through $N_{s}\sim t_{f}\sim\sqrt{\omega}$, the energy reduction prescription gives a larger contribution. This has two origins. On the one side, the energy loss in the individual elastic collisions (see Fig. 28) drives some gluons to the soft sector for which the formation time (and so the associated $N_{s}$) is smaller. Because for finite gluon masses the formation time is $\propto\omega/(m_{g}^{2}+k_{{\perp}}^{2})$, the shift of the gluons to smaller $k_{{\perp}}$ values, seen in Fig. 29, does not lower the formation time considerably. For the ‘‘energy reduction’’ scenario we observe therefore an increase of the production for low $N_{S}$. This effect is reinforced by a kinematic effect: The smaller the value of $\omega$ the larger is the probability that elastic collisions of the virtual gluons are kinematically not possible. Then the collisions do not take place – as already described above – and $N_{s}$ is not increased but the phase accumulation continues. This also leads to an enhancement of the entry at $N_{s}=1$, as seen in Fig. 30.
@@ V.4 Testing the role of the phase accumulation
Another quantity, which is approximated in the BDMPS-Z conditions, is the phase accumulation of the virtual gluon due to elastic collisions. Figs. 32, 32, and 33 show the influence of different forms of the phase accumulation of the gluon emission amplitudes on the gluon observables. The form of the phase accumulation cannot be inferred from more fundamental approaches. As discussed in section IV the virtual gluon will be considered as a formed particle once its phase passes the critical value of $\varphi_{c}=6$. We display the results for three different descriptions of this phase accumulation. The results for the phase increment of $\Delta\varphi=\frac{2P\cdot k^{\prime}}{E}\cdot\Delta t$ are shown as a blue solid line, that of a phase increment of $\Delta\varphi=\frac{k_{\perp}^{2}}{\omega}\cdot\Delta t$, corresponding to a vanishing gluon mass, as a dashed orange line and that of a phase increment of $\Delta\varphi=\frac{m_{g}^{2}+k_{\perp}^{2}}{\omega}\cdot\Delta t$, corresponding to the limit $E\to\infty$, as a green dotted line.
@@
For both, the gluon $\omega$ and the $k_{\perp}$ spectra, the high energy limit $\Delta\varphi=\frac{m_{g}^{2}+k_{\perp}^{2}}{\omega}\cdot\Delta t$ presents a reasonable approximation of the more general formula $\Delta\varphi=\frac{2P\cdot k^{\prime}}{E}\cdot\Delta t$.